OpenStudy (anonymous):

If f(x)=2x^2−4x−3 and g(x)=(x−4)(x+4), find f(x)−g(x).

OpenStudy (anonymous):

4x+13 x2−4x+13 3x2−4x+13 −4x2−4x+13

OpenStudy (anonymous):

g(x) = x^2 - 16 using identity (a-b)(a+b) = a^2 - b^2

OpenStudy (anonymous):

Alright I can help. So Im gonna give you an example

OpenStudy (anonymous):

so f(x) - g(x) =2x^2−4x−3 - (x^2 - 16) = x^2 - 4x + 13

OpenStudy (anonymous):

still so confused

OpenStudy (anonymous):

please help, i have to finish my class by 9 pm today

OpenStudy (anonymous):

f(x) - g(x) = 2x^2−4x−3 - (x^2 - 16) = 2x^2 - 4x - 3 - x^2 + 16 = 2x^2 - x^2 - 4x - 3 + 16 = x^2 - 4x + 13 is it clear now??

OpenStudy (anonymous):

kinda not really

OpenStudy (anonymous):

Is this part clear??? g(x) = x^2 - 16 using identity (a-b)(a+b) = a^2 - b^2